# ComGPT: Detecting Local Community Structure with Large Language Models

### v1

@@ Abstract
Large language models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks such as node classification. So far, how to leverage LLMs to better detect local communities remains underexplored. Local community detection algorithms based on seed expansion often face a seed-dependent problem, community diffusion, and free rider effect. Using LLMs to solve existing local community work problems faces the following challenges: existing graph encoding methods fail to provide LLMs with sufficient community-related graph information; LLMs lack domain knowledge in mining communities. To address these issues, we improve graph encoding by supplementing community knowledge to enhance the ability of graph encoding to express graph information. Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics, aiming to alleviate the seed-dependent problem, community diffusion, and free rider effect. Based on the graph encoding and NSG prompt, we present a GPT-guided local community detection, called ComGPT. ComGPT iteratively selects potential nodes from the detected community’s neighbors and subsequently employs GPT to choose the node that optimally integrates into the detected community from these selected potential nodes. Experimental results demonstrate that ComGPT outperforms the comparison algorithms, thereby confirming the effectiveness of the designed graph encoding and prompt.
@@ Introduction
Identifying community structures from networks has wide applications, such as identifying protein complexes (Yin et al. 2017). Local community detection has gained attention for its ability to quickly identify the community containing a specific node (Ni et al. 2024a, b). Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), $k$-core decomposition (Cui et al. 2014), $k$-truss identification (Faysal et al. 2023), and non-negative matrix factorization algorithms (Wang et al. 2011). Among these methods, the seed expansion algorithm has garnered widespread acclaim for its versatility and ease of implementation. The seed expansion algorithms typically enlarge the community by selecting a node with the highest quality function value at each step. However, it often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address these problems, researchers developed different strategies such as identifying core node sets and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018).
@@
In this study, we explore an alternative approach inspired by the the success of large language models (LLMs) in many fields (Gu et al. 2024; Zong and Krishnamachari 2024; Yang et al. 2022) such as theory-of-mind reasoning (Adhikari et al. 2020; Ammanabrolu and Riedl 2021) and graphical reasoning (Wang et al. 2023). As far as we know, LLMs have not been applied to local community detection. Therefore, we explore the application of large models for local community detection and propose an algorithm, called ComGPT (GPT-guided Local Community Detection). This algorithm leverages GPT to guide the detection process, thereby bridging the current gap in LLMs applicability to this field. ComGPT integrates the seed expansion algorithm with GPT, leveraging their inherent compatibility: GPT through step-by-step guidance, while the seed expansion method incrementally adds nodes to the community, making their integration highly synergistic.
@@
However, the application of LLMs to local community detection faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information. To address this, we enhance graph encoding by incorporating two key aspects: graph topology and community information. 2) Lacks domain knowledge for LLMs in dealing with local community detection. It limits the capabilities of LLMs. To address this, we have designed prompt to assist LLMs in understanding community definitions and how to select nodes. 3) Lacks efficient interaction between GPT and community detection methods. The mode for integrating community detection algorithms with GPT, as well as strategies for reducing the number of input tokens during their interaction, remain unexplored. Based on the step-by-step characteristics of GPT and seed expansion algorithm, we design a step-by-step interactive mode. To minimize costs by reducing the number of tokens, the input to GPT is not the entire graph but a partial graph that includes the detected community and some of its neighbors. Our contribution is summarized as follows:
@@
  • We propose a GPT-guided local community detection algorithm, called ComGPT. ComGPT repeatedly selects potential nodes from the neighbors of the detected community and then uses GPT to identify the node that optimally integrates into the detected community from these selected potential nodes. To the best of our knowledge, this is the first work to apply LLMs to the domain of local community detection.
@@
  • For GPT to understand and detect communities, we design graph encoding method, NSG prompt, and interaction mode. The graph encoding method incorporates both graph topology and community information, which enhances GPT to understand the current state of the detected community. NSG prompt that integrates community features facilitates GPT’s comprehension of domain knowledge. Additionally, an interactive mode is established between GPT and the local community detection method.
@@
  • The results on real datasets indicate that ComGPT outperforms comparison methods. Ablation study confirm that GPT plays a crucial role in the ComGPT ’s performance. Furthermore, prompt comparisons experiments reveal that the designed prompt effectively aids GPT in understanding the graph structure and community situation.
@@
Local community detection focuses on discovering the community to which the seed node belongs (Ni et al. 2024c, 2023). A variety of methods like seed expansion algorithms are developed for detecting local communities (Luo, Wang, and Promislow 2008; Cui et al. 2014), which expand communities by selecting nodes with the highest quality function values. For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the highest local modularity R and M, respectively. The seed expansion algorithms often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al. 2022), finding the nearest nodes with greater centrality (Luo et al. 2020), and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018). Our work mitigates the aforementioned problems by utilizing GPT, distinguishing it from existing studies.
LLMs
@@
In recent years, researchers have explored LLMs for node classification (Chen et al. 2023) and graph reasoning (Zhang 2023). Wang et al. demonstrated LLMs’ graph reasoning abilities and designed the NLGraph dataset (Wang et al. 2023) to evaluate these capabilities. Furthermore, Guo et al. emphasized the design and improvement of graph encoding techniques, which critically affect GPT’s interpretation of input data and its subsequent output (Guo et al. 2023). Therefore, Adjacency and Incident (Fatemi, Halcrow, and Perozzi 2023) are designed to encode graph structured data into text for LLMs. Adjacency uses integer node encoding and parenthesis edge encoding. Incident uses integer node encoding and incident edge encoding. Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection. Our work introduces a novel graph encoding method tailored for the community detection domain, which incorporates community structure considerations, distinguishing it from existing approaches.
@@
Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al. 2024), such as the general COT prompt (Wei et al. 2022) and the BAG prompt (Wang et al. 2023) for graph reasoning tasks. However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks. This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above work.
Proposed algorithm
Shortcomings
Existing heuristic algorithms suffer from the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al. 2013), which are explained as following.
(a) Seed-dependent problem
(b) Community diffusion
(c) Free rider effect
Figure 1: Existing shortcomings. Nodes in the same community are marked with the same color.
Shortcoming 1: Seed-dependent problem (Kloumann and Kleinberg 2014).
It refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(1(a)). The communities detected by the M method (Luo, Wang, and Promislow 2008) with nodes 5 and 1 as the starting seed nodes are [5, 11, 6, 7, 17] and [1, 12, 13, 4, 8, 2, 14, 3, 22, 18, 20], respectively. The former includes some nodes of the real community and misses many nodes, while the latter is more accurate.
Shortcoming 2: Community diffusion.
The community diffusion problem means that the detected community contains nodes from different communities. Take node 1 in Figure 1(1(b)) as an example to illustrate the community diffusion problem. The community detected by the M method with node 1 as the starting seed node is [1, 2, 3, 4, 7, 8, 9, 13, 14, 15, 16, 17, 18, 20], which includes nodes from two different communities.
Shortcoming 3: Free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al. 2013).
The free rider effect refers to mistakenly adding unrelated communities to enhance the fitness of a detected community. We use the community derived from seed node 7, as shown in Figure 1(1(c)), to illustrate it. The target community consists of nodes 5, 6, 7, 8, and 9. However, to achieve a higher fitness value, the detected community not only contains nodes in the target community [5, 6, 7, 8, 9] but also erroneously includes all nodes in the unrelated community [1, 2, 3, 4, 10, 11].
ComGPT
Figure 2: The diagram of ComGPT.
@@
We propose a local community detection algorithm based on GPT, called ComGPT. Figure 2 shows the main steps of ComGPT. It includes four main steps: 1) potential nodes identification; 2) graph encoding; 3) node selection by GPT; 4) node supplementation by GPT. Potential nodes identification aims to pinpoint potential nodes from the neighbors of the current community. Graph encoding involves transforming the graph structure into a textual format interpretable by GPT. Node selection is that GPT selects the node that optimally integrates into the community from potential nodes. Node supplement involves GPT selecting which nodes to add to the community when no potential nodes remain.
@@
Input: Network $G$, Seed node $v_{s}$
@@
5:        $gtxt\leftarrow$ graph encoding with $C$ and $N$
6:        $v\leftarrow$ node selection by GPT from $PN$
@@
10:     $v\leftarrow$ node supplementation by GPT
11:     if $v$ is not null then
12:        add $v$ to $C$
13:     else
14:        break
15:     end if
16:  end while
17:  $C\leftarrow$ community with highest $M$ in $cands$
18:  return  $C$
Algorithm 1 shows the overall process of ComGPT. For convenience, we briefly introduce the notations used in this paper. Let $G$ denote the network, $v_{s}$ denote the seed node, $C$ denote the detected community, $v$ denote the node returned by the node selection or node appending steps, $gtxt$ denote the text of the graph structure obtained through graph encoding, $PN$ denote the set of potential nodes, $cands$ store the community each time $PN$ is empty. Initially, community $C$ consists of only one seed node $v_{s}$ and the $cands$ is empty (line 1). First, ComGPT selects potential nodes $PN$ from neighbors of current community $C$ for GPT (line 3). The local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$, is converted into text, i.e., $gtxt$ (line 5). Utilizing $gtxt$ and NSG (Node Select Guide) prompt, GPT selects the node $v$ from potential nodes $PN$ and adds it to the current community $C$ (line 6,7). The above steps are repeated until $PN$ is empty. After the potential nodes $PN$ is empty, the community $C$ is added to the candidate communities $cands$ (line 9). When $PN$ is empty, it does not necessarily imply that the neighboring nodes of the community are weakly connected to the community. Instead, it could be that the nodes have relatively high degrees, which marginally reduces the $M$ value. To address this scenario, ComGPT identifies nodes through node supplementation (line 10). If GPT identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 12). Otherwise, ComGPT stops community expansion (line 14). After the above process is completed, ComGPT selects the community with the highest $M$ value from $cands$ as the final community (line 17).
@@
To mitigate the high computational cost associated with processing nodes through GPT, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT for selection. Different fitness functions are designed in the seed expansion algorithm to select nodes, which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al. 2023). ComGPT uses local modularity $M_{C}$ to select potential nodes, which is defined as the ratio of the number of internal edges in community $C$ to the number of external edges of community $C$. Nodes that are satisfied $\Delta M=M_{C\cup\{v\}}-M_{C}>0$ are regarded as potential nodes. During the initial expansion of the community, since nodes with high degrees often struggle to achieve $\Delta M>0$, these nodes are not considered potential nodes $PN$. To alleviate this preference, these nodes are added to potential nodes based on their closeness value to the current community $C$. The closeness value $f_{v}$ between node $v$ and the community $C$ is calculated as
$f_{v}=\frac{|C\cap N(v)|}{|C|},$ (1)
$|C|$ represents the number of nodes in the community and $N(v)$ represents the neighbors of node $v$. The value of $f_{v}$ ranges from 0 to 1. Higher values indicate a greater proportion of nodes connected to those within the community.
The initial potential nodes consist of nodes with $\Delta M>0$ and nodes with high $f_{v}$ values. To further reduce the number of nodes processed by GPT, $k$ nodes are selected from them to form a potential node’s set.
@@
Figure 3: Graph encoding.
@@
The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT. The graph data includes two aspects. 1) Graph topological structure. We adopt the Incident method (Fatemi, Halcrow, and Perozzi 2023) to let GPT understand the graph’s topological structure. The Incident method represents the graph topology by introducing the first-order neighbors of each node. To reduce computational overhead, the graph structure inputted to GPT is a local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$. For clarity, the middle part of Figure 3 illustrates the specific graph encoding: “G describes a graph among nodes:1, 2, 3, 5, 6, 7. Node 1 is connected to nodes 2,3”. To reduce the number of tokens, some conjunctions that do not affect the understanding of GPT are removed. 2) Supplementary knowledge. We describe node information from a community perspective to enhance GPT’s understanding of community structure. This involves outlining which nodes are within the community, those that are outside, and detailing their specific external connections. For clarity, the lower part of Figure 3 illustrates graph encoding for community $C$ = [1, 2, 3] and potential nodes $PN$ = [5, 6, 7]: “Nodes in the current community:[1,2,3]. The outside nodes contain:[5,6,7]. Node 5 is connected to nodes in community:[1,3]”.
Node selection by GPT
Node selection refers to choosing the most suitable node for community expansion, based on the prompt and graph data. Below, we introduce node selection through prompts that teach GPT to accurately select nodes.
@@
1. The more a potential node is connected to other potential nodes, the higher the likelihood of its selection: This prompt could alleviate the seed-dependent problem and the free rider effect.
@@
The reason for the seed-dependent problem is low-degree node preference. Take seed node 5 in Figure 1(1(a)) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it. Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges. Thus, the M method selects node 11 to join the community. Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem. Taking Figure 1(1(a)) as an example, the community includes node 5, with nodes 1, 7, and 11 regarded as potential nodes. According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
@@
The reason for the free rider effect is the neglect of the connectivity of neighbor nodes of the detected community. Nodes that connect to many other neighbor nodes are consideration of their association with other potential nodes, which alleviates the free rider effect. Taking Figure 1(1(c)) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes. According to prompt 1, nodes 6 and 8 have a higher probability of being selected. Node 4 is a node that may cause the free rider effect, and will not be selected.
@@
2. Prioritize selecting nodes that are connected to multiple nodes within the community: This prompt could alleviate the community diffusion problem. The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community. Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem. Taking Figure 1(1(b)) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where node 13 or 18 may lead to community diffusion. According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
The first prompt is to consider the closeness with neighboring nodes outside the community, and the second prompt is to consider the closeness with nodes within the community. This is consistent with the characteristics of the community being close inside and sparse outside, thus helping GPT understand the characteristics of the nodes to be selected.
Node supplementation by GPT
@@
When $PN$ is empty, ComGPT uses GPT to supplement nodes. Node supplementation involves evaluating whether to continue the expansion process or terminate it. If continuation is warranted, it also determines which additional nodes should be selected. In terms of graph encoding, the graph encoding method mentioned above, including graph topological structure and supplementary knowledge, is used. The difference is that to provide GPT with sufficient information to make a decision, the input graph text corresponds to graph topology that includes the current community and its first and second-order neighbors. The nodes to be selected by GPT are the three nodes with the largest M values among the first-order neighbor nodes of the community. In terms of prompt, it is based on the definition of the community and whether these nodes can maintain the cohesion of the community. If GPT determines that any of these three nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns NULL.
@@
To evaluate the performance of our algorithm, we conducted some experiments on real datasets. We first outline the experimental settings, which include datasets, comparison algorithms, and evaluation metrics. We then present the experimental results and provide corresponding analyses.

### v2

@@ Abstract
Large Language Models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks, such as node classification. Despite their strong abilities in graph reasoning tasks, they lack specific domain knowledge and have a weaker understanding of community-related graph information, which hinders their capabilities in the community detection task. Moreover, local community detection algorithms based on seed expansion, referred to as seed expansion algorithms, often face the seed-dependent problem, community diffusion, and free rider effect. To use LLMs to overcome the above shortcomings, we explore a GPT-guided seed expansion algorithm named ComGPT. ComGPT iteratively selects potential nodes by local modularity $M$ from the detected community’s neighbors, and subsequently employs LLMs to choose the node to join the detected community from these selected potential nodes. To address the above issues faced by LLMs, we improve graph encoding method, called Incident, by incorporating community knowledge to improve LLMs’s understanding of community-related graph information. Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics. Experimental results demonstrate that ComGPT outperforms the comparison methods, thereby confirming the effectiveness of the improved graph encoding method and prompts.
@@ Introduction
Identifying community structures from networks has wide applications, such as identifying protein complexes (Yin et al. 2017). Local community detection has gained attention due to its ability to quickly identify the community containing a specific node (Ni et al. 2024a, b). Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), and non-negative matrix factorization algorithms (Wang et al. 2011). Among these methods, the seed expansion algorithms have garnered widespread acclaim for their versatility and ease of implementation. The seed expansion algorithms typically enlarge the community by selecting a node with the high scoring function value at each step (Moradi, Olovsson, and Tsigas 2014).
However, seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013), shown in Figure 1. The seed-dependent problem refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(a). The community diffusion problem means that the detected community contains nodes from different communities, as illustrated by node 1 in Figure 1(b). The free rider effect refers to mistakenly adding unrelated communities to enhance the scoring function value of a detected community, as illustrated by node 7 in Figure 1(c). To address these problems, researchers developed different strategies such as identifying core node sets (Guo et al. 2022) and designing scoring functions with a bias towards nodes with large degrees (Luo et al. 2018).
((a))
((b))
((c))
Figure 1: Some shortcomings faced by seed expansion algorithms. Nodes in the same community are marked with the same color.
@@
In this study, inspired by the success of Large Language Models (LLMs) in many fields (Gu et al. 2024; Zong and Krishnamachari 2024; Yang et al. 2022)(Adhikari et al. 2020; Ammanabrolu and Riedl 2021)(Wang et al. 2023), we explore an alternative approach based on LLMs, called ComGPT (GPT-guided seed expansion algorithm). ComGPT leverages LLMs to guide the local community detection process to alleviate the seed-dependent problem, community diffusion, and free rider effect. The one-question-one-answer modality of LLMs aligns well with the incremental nature of seed expansion algorithms. It bridges the current gap in LLMs applicability to the community detection domain.
@@
Specifically, we leverage LLMs to select nodes to join the community instead of the scoring function commonly used in seed expansion algorithms. However, the application of LLMs to select nodes faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information. To address this, we improve the graph encoding method, called Incident (Fatemi, Halcrow, and Perozzi 2023), by incorporating community information. 2) Lack of domain knowledge for LLMs in dealing with local community detection. It limits the capabilities of LLMs. To address this, we design prompts to assist LLMs in understanding community definitions and how to select nodes. Additionally, LLMs receive only a partial network as input, not the entire network, to minimize costs by reducing the number of tokens processed.
Our contributions are summarized as follows:
@@
  • We propose a GPT-guided seed expansion algorithm called ComGPT, which iteratively uses GPT-3.5-turbo to add nodes to the community. It includes four main steps: potential node identification, graph encoding, node selection by GPT-3.5-turbo, and node supplementation by GPT-3.5-turbo. To the best of our knowledge, this is the first work to apply LLMs to the domain of community detection.
@@
  • We design an improved version of the typical graph encoding method, named Incident (Fatemi, Halcrow, and Perozzi 2023), and NSG (Node Selection Guide) prompt. This improved version incorporates both graph topology and community information, which enhances LLMs to understand the current state of the detected community. NSG prompt that integrates community features facilitates LLMs’s comprehension of domain knowledge.
@@
  • The results on real datasets indicate that ComGPT outperforms comparison methods. The ablation study experiments confirm that LLMs play a crucial role in the ComGPT’s performance. Furthermore, prompt comparison experiments reveal that the designed prompt effectively aids LLMs in understanding the graph structure and community situation.
@@
Local community detection focuses on discovering the community to which the seed node belongs (Ni et al. 2024c, 2023). A variety of methods, like $k$-core decomposition (Cui et al. 2014) and seed expansion algorithms (Clauset 2005), are developed for detecting local communities. Seed expansion algorithms expand communities by selecting nodes with high scoring function values (Moradi, Olovsson, and Tsigas 2014). For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the local modularity R and M, respectively. The seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al. 2022; Ji, Guo, and Yu 2022), finding the nearest nodes with greater centrality (Luo et al. 2020), alternating strategy of strong fusion and weak fusion (Shang et al. 2023), and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018). Our work mitigates the aforementioned problems by utilizing LLMs, distinguishing them from existing studies.
LLMs in graph reasoning
@@
In recent years, researchers have explored the application of LLMs in graph reasoning (Zhang 2023; Chen et al. 2023). Wang et al. demonstrated LLMs’ graph reasoning abilities and designed the NLGraph dataset (Wang et al. 2023) to evaluate these capabilities. Furthermore, Guo et al. emphasized the design and improvement of graph encoding techniques, which critically affect LLMs’ interpretation of input data and its subsequent output (Guo et al. 2023). Therefore, Adjacency and Incident (Fatemi, Halcrow, and Perozzi 2023) are designed to encode graph structured data into text for LLMs. Adjacency uses integer node encoding and parenthesis edge encoding. Incident uses integer node encoding and incident edge encoding. Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection. Our work introduces a novel graph encoding method based on Incident (Fatemi, Halcrow, and Perozzi 2023), tailored for the community detection domain, that incorporates community structure and distinguishes itself from existing approaches.
@@
Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al. 2024), such as the general CoT (Chain-of-Thought) prompt (Wei et al. 2022) and the BaG (Build a Graph) prompt (Wang et al. 2023) for graph reasoning tasks. However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks. This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above prompts.
Approach
@@
Figure 2: The diagram of ComGPT. It includes four main steps: 1) Potential node identification aims to pinpoint potential nodes from the neighbors of the current community. 2) Graph encoding involves transforming the graph structure into a textual format interpretable by GPT-3.5-turbo. 3) Node selection by GPT-3.5-turbo is that GPT-3.5-turbo selects the node that optimally integrates into the community from potential nodes. 4) Node supplementation by GPT-3.5-turbo involves GPT-3.5-turbo selecting which nodes to add to the community when no potential nodes remain.
We introduce a GPT-guided seed expansion algorithm named ComGPT, based on GPT-3.5-turbo. This section begins with an overview of ComGPT, followed by detailed descriptions of each step of ComGPT.
Overview
@@
Input: Network $G$, Seed node $v_{s}$, Prompts NSG and NSU for node selection and supplementation, respectively, Questions Qsel and Qsup for node selection and supplementation, respectively
@@
5:        $gtxt\leftarrow$ graph encoding with ($C\cup PN$, $C$, $PN$)
6:        $v\leftarrow$ node selection by GPT-3.5-turbo with ($gtxt$, NSG, Qsel )
@@
10:     $N\leftarrow$ first and second-order neighbors of $C$
11:     $PN\leftarrow$ nodes selected by modularity $M$
12:     $gtxt\leftarrow$ graph encoding with ($C\cup N$, $C$, $PN$)
13:     $v\leftarrow$ node supplementation by GPT-3.5-turbo with ($gtxt$, NSU, Qsup)
14:     if $v$ is not null then
15:        add $v$ to $C$
16:     else
17:        break
18:     end if
19:  end while
20:  $C\leftarrow$ community with highest $M$ in $cands$
21:  return  $C$
As shown in Figure 2, ComGPT contains four main steps. To better understand the interactions among these steps, Algorithm 1 shows the overall process of ComGPT. Initially, community $C$ contains seed node $v_{s}$ and $cands$ is empty (line 1), where $cands$ stores the community whenever the potential node identification step returns an empty set. First, ComGPT selects potential nodes $PN$ with the step of potential node identification (line 3), used for subsequent graph encoding. In order for GPT-3.5-turbo to understand the graph structure and community information, the local network, i.e., the subgraph consisting of the nodes in both current community $C$ and $PN$ as well as edges between these nodes, is converted into text, i.e., $gtxt$ (line 5). Utilizing $gtxt$, NSG (Node Select Guide) prompt, and Qsel, GPT-3.5-turbo selects node $v$ from potential nodes $PN$ and adds it to the current community $C$ (lines 6-7). The above steps are repeated until $PN$ is empty. If $PN$ is empty, the community $C$ is added to the $cands$ (line 9). The fact that $PN$ is empty does not necessarily imply that the neighboring nodes of the community are weakly connected to the community. Therefore, based on $gtxt$, NSU (Node SUpplementatio) prompt, and Qsup, ComGPT continues to identify nodes through node supplementation (lines 10-13). If GPT-3.5-turbo identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 15). Otherwise, GPT-3.5-turbo returns “null” and ComGPT stops community expansion (line 17). After the above process is completed, ComGPT selects the community with the highest local modularity $M$ calculated by formula (1) from $cands$ as the final community (line 20).
@@
To mitigate the cost associated with processing nodes through GPT-3.5-turbo, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT-3.5-turbo for selection. Different scoring functions are designed in the seed expansion algorithms to select nodes (Moradi, Olovsson, and Tsigas 2014), which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al. 2023). ComGPT uses local modularity $M_{C}$ (Luo, Wang, and Promislow 2008) to select potential nodes. $M_{C}$ is calculated as,
$M_{C}=\frac{e_{ic}}{e_{oc}},$ (1)
where $e_{ic}$ represents the number of internal edges in community $C$ and $e_{oc}$ represents the number of external edges of community $C$. For each first-order neighbor node $v$ of community $C$, if node $v$ satisfies $\Delta M=M_{C\cup\{v\}}-M_{C}>0$, then it is considered a potential node. To further reduce the number of nodes processed by GPT-3.5-turbo, $k$ nodes with the largest $M$ are selected from the above potential nodes to form a potential node’s set $PN$.
@@
Figure 3: An toy example of graph encoding. The middle part illustrates the text corresponding to graph topological structure. The lower part illustrates the text corresponding to supplementary community information.
@@
The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT-3.5-turbo. The graph data includes two aspects. 1) Graph topological structure. We adopt Incident (Fatemi, Halcrow, and Perozzi 2023) to let GPT-3.5-turbo understand the graph’s topological structure. Incident represents the graph topology by introducing the first-order neighbors of each node. To reduce computational overhead, the graph structure input to GPT-3.5-turbo is a local network. For the node selection by GPT-3.5-turbo step, the local network comprises nodes in $C\cup PN$ and the edges between them. For the node supplementation by GPT-3.5-turbo step, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$. 2) Supplementary knowledge. We describe node information from a community perspective to enhance GPT-3.5-turbo’s understanding of community structure. This step involves outlining which nodes are within the community, those that are outside, and the connections of outside nodes to both within and outside the community. The outside nodes are potential nodes in $PN$ from the step of potential nodes identification. The text corresponding to the topological structure and the text corresponding to the supplementary information are concatenated to form $gtxt$. To better understanding, an example is given in Figure 3.
Node selection by GPT-3.5-turbo
Node selection refers to choosing the most suitable node for community expansion, based on NSG prompt, graph text $gtxt$, and question Qsel. Graph text $gtxt$ is obtained by the step of graph encoding with the local network, $C$, and $PN$, where the local network consists of the nodes in $C\cup PN$ and edges between these nodes. Below, we introduce NSG prompt and question Qsel.
@@
  • Prompt 1: The more an outside node is connected to other outside nodes, the higher the likelihood of its selection. The outside nodes provided to GPT-3.5-turbo are potential nodes from the step of potential nodes identification. This prompt could alleviate the seed-dependent problem and the free rider effect.
@@
    The reason for the seed-dependent problem is low-degree node preference. Take seed node 5 in Figure 1(a) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it. Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges. Thus, the M method (Luo, Wang, and Promislow 2008) selects node 11 to join the community. Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem. Nodes 1, 7, and 11 are potential nodes for community [5]. According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
@@
    The reason for the free rider effect is adding nodes that are sparsely connected to the community to the community. Nodes that are connected to multiple potential nodes indirectly maintain close connections to the community, which alleviates the free rider effect. Taking Figure 1(c) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes. According to prompt 1, nodes 6 and 8 have a higher probability of being selected. Node 4 is a node that may cause the free rider effect, and will not be selected.
@@
  • Prompt 2: Prioritize selecting outside nodes that are connected to multiple nodes within the community. The outside nodes mentioned in prompts are potential nodes from the step of potential nodes identification. This prompt could alleviate the community diffusion problem. The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community. Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem. Taking Figure 1(b) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where nodes 13 or 18 may lead to community diffusion. According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
    The first prompt is to consider the closeness of the node to potential nodes outside the community, and the second prompt is to consider the closeness of the node to nodes within the community. This is consistent with the characteristics of the community being close inside, thus helping GPT-3.5-turbo understand the characteristics of the nodes to be selected. To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for selecting nodes is designed: “You’re doing local community detection. Based on the graph data and prompt, please select a node that you think is most likely to belong to the current community $C$ for community expansion. Provide a detailed explanation.”. Based on the above prompts, the text $gtxt$ obtained by graph encoding method, and the question Qsel, GPT selects the node $v$ from potential nodes.
Node supplementation by GPT-3.5-turbo
@@
Node supplementation involves evaluating whether to continue the expansion process or terminate it. If continuing, it also determines which additional nodes should be selected. Next, we introduce the information needed for graph encoding, NSU prompt, and question Qsup in turn. For graph encoding, to provide GPT-3.5-turbo with sufficient information to make a decision, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$. Besides, ComGPT forcibly adds the $k$ nodes with the highest $M$ values from the first neighbors of community $C$ to $PN$. NSU prompt is based on the definition of the community and whether these nodes can maintain the cohesion of the community. To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for supplementing nodes is designed: “Please analyze whether these nodes should be added to the community $C$. The probability of not adding nodes is higher. But it doesn’t mean you always refuse to add nodes. If you think there is a suitable node, please output its node number.”. Based on the above prompts, the text obtained by graph encoding method, and the question Qsel, if GPT-3.5-turbo determines that any of nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns “null”.